narutouzamaki2:

A quarter has a diameter of approximately 24 millimeters. What is its approximate circumference? 75.36 mm 150.72 mm 7.64 mm 37.68 mm

narutouzamaki2:

plz help

narutouzamaki2:

help me

narutouzamaki2:

r yall gonna help me or not

ImBadKidTae:

what do you know

ImBadKidTae:

: Circumference = 75.36 mm. Step-by-step explanation: Given : A quarter has a diameter of approximately 24 millimeters.

jhonyy9:

@XioGonz this quarter mean here a quarter of a sphere - why not specified ?

XioGonz:

The approximate Circumference of the quarter is 75.398mm. Why? C=Dπ, D=2, and C=75.398mm

XioGonz:



jhonyy9:

so what mean here the quarter ?

ImBadKidTae:

ok

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so what mean here the quarter ? \(\color{#0cbb34}{\text{End of Quote}}\) An actual quarter

jhonyy9:

quarter is 1/4 or 25% but of what ?

jhonyy9:

ohh of dollar cents quarter ?

XioGonz:

Yes

jhonyy9:

so than this is very easy diameter 24 mm => r = d/2 = ? circumference formula = 2*pi*r = 2*3,14*r hope helped

jhonyy9:

do you think it so ?

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 do you think it so ? \(\color{#0cbb34}{\text{End of Quote}}\) You are correct, I agree!

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 do you think it so ? \(\color{#0cbb34}{\text{End of Quote}}\) You are correct, I agree! \(\color{#0cbb34}{\text{End of Quote}}\) ty

ImBadKidTae:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 do you think it so ? pew pew pew \(\color{#0cbb34}{\text{End of Quote}}\) ty \(\color{#0cbb34}{\text{End of Quote}}\)???